a shell fired from ground is just able to cross horizontally the top of a wall 90 m away and 45 of height.find the angle of projection of shell with the horizontal . 
(ans is 45degree)

Grade:11

1 Answers

Srijan
6 years ago
Let horizontal velo. be V. Very velo. be v.v^2=2gh =2×9.8×45 v=21t=21/9.8V=90/t =42Angle= tan inverse (v/V) =tan^-1(1/2) =45°